The mean age of the 6 women in a class is 23. The mean age of the 10 men in the same class is 26. What is the mean age of the 16 students in the class?
Found 2 solutions by mananth, vksarvepalli:
Answer by mananth(16949)

The mean age of the 6 women in a class is 23. The mean age of the 10 men in the same class is 26. What is the mean age of the 16 students in the class?
the total age of all women = 6*23 =138
the total age of all men = 10*26 =260
the total age of all students = 398
the number of students = 16
average age in the class = 398/16


Answer by vksarvepalli(154)

The mean age of the 6 women in a class is 23
so (sum of all ages of women)/6 = 23
so sum of ages of all 6 women = 23*6 = 138
The mean age of the 10 men in the same class is 26
so sum of ages of all 10 men = 26*10 = 260
sum of ages of all the 16 people= 138+260 = 398
so mean age of the class = 398/16 = 24.875